# A new framework for nonlinear Kalman filters

### v10

@@ 2.3 The proposed new framework for nonlinear Kalman filters
Note that the proposed new framework selects the same Kalman gain and does the same state update as the conventional one; the only difference is that the state covariance update equation is now a direct approximation of (11), which is the actual state covariance regardless of the selection of the Kalman gain $K$. Intuitively, this can prevent the KF from overconfidence because the equation no longer relies on the problematic assumption that the Kalman gain is optimal. To explain this advantage more rigorously, we consider the newly approximated covariance matrices ${\color[rgb]{1,0,0}P_{y,k|k}},{\color[rgb]{1,0,0}S_{k|k}}$, and ${\color[rgb]{1,0,0}P_{xy,k|k}}$ as noisy versions of $P_{y,k},S_{k}$, and $P_{xy,k}$. Similar to our previous assumptions about ${\color[rgb]{1,0,0}P_{y,k|k-1}},{\color[rgb]{1,0,0}S_{k|k-1}}$, and ${\color[rgb]{1,0,0}P_{xy,k|k-1}}$, we assume that ${\color[rgb]{1,0,0}P_{xy,k|k}}$ is a random matrix, ${\color[rgb]{1,0,0}P_{y,k|k}}$ is a random positive semidefinite matrix, and ${\color[rgb]{1,0,0}S_{k|k}}$ is a random positive definite matrix. Additionally, we assume that the “noises” that arise in these approximations are unbiased, so these new random matrices should also satisfy:
@@
$\begin{cases}\operatorname{\mathbb{E}}[{\color[rgb]{1,0,0}P_{y,k|k}}]=P_{y,k}\\ \operatorname{\mathbb{E}}[{\color[rgb]{1,0,0}S_{k|k}}]=S_{k}\\ \operatorname{\mathbb{E}}[{\color[rgb]{1,0,0}P_{xy,k|k}}]=P_{xy,k}\end{cases}$ (19)
@@
Assume that $\operatorname{\mathbb{E}}({\color[rgb]{1,0,0}P_{xy,k|k-1}})=\operatorname{\mathbb{E}}({\color[rgb]{1,0,0}P_{xy,k|k}})=P_{xy,k}$ and $\operatorname{\mathbb{E}}({\color[rgb]{1,0,0}S_{k|k-1}})=\operatorname{\mathbb{E}}({\color[rgb]{1,0,0}S_{k|k}})=S_{k}$. If ${\color[rgb]{1,0,0}P_{xy,k|k-1}}$, ${\color[rgb]{1,0,0}P_{xy,k|k}}$, ${\color[rgb]{1,0,0}S_{k|k-1}}$, and ${\color[rgb]{1,0,0}S_{k|k}}$ are all independent from each other, ${\color[rgb]{1,0,0}P_{k|k,new}}$ calculated in (18) is an unbiased estimation of ${\color[rgb]{1,0,0}P_{k|k,ac}}$. Namely, we have:
$\operatorname{\mathbb{E}}({\color[rgb]{1,0,0}P_{k|k,new}}-{\color[rgb]{1,0,0}P_{k|k,ac}})=0$ (20)
@@
$\operatorname{\mathbb{E}}({\color[rgb]{1,0,0}P_{k|k,new}}-{\color[rgb]{1,0,0}P_{k|k,ac}})$ (21)
$=$ $\operatorname{\mathbb{E}}[{\color[rgb]{1,0,0}P_{xy,k|k-1}}{\color[rgb]{1,0,0}S_{k|k-1}^{-1}}({\color[rgb]{1,0,0}S_{k|k}}-S_{k}){\color[rgb]{1,0,0}S_{k|k-1}^{-1}}{\color[rgb]{1,0,0}P_{xy,k|k-1}^{T}}]$
$-\operatorname{\mathbb{E}}[({\color[rgb]{1,0,0}P_{xy,k|k}}-P_{xy,k}){\color[rgb]{1,0,0}S_{k|k-1}^{-1}}{\color[rgb]{1,0,0}P_{xy,k|k-1}^{T}}]$
$-\operatorname{\mathbb{E}}[{\color[rgb]{1,0,0}P_{xy,k|k-1}}{\color[rgb]{1,0,0}S_{k|k-1}^{-1}}({\color[rgb]{1,0,0}P_{xy,k|k}}-P_{xy,k})^{T}]$
$=$ 00 0
@@
The last equality equals zero because the terms in parentheses are zero, due to the unbiased (e.g. $\operatorname{\mathbb{E}}({\color[rgb]{1,0,0}S_{k|k}})=S_{k}$) and independence assumptions.
@@
Theorem 2 suggests that the proposed new framework can generally prevent the filter from being overconfident if the approximation errors of the measurement function at the predicted and updated states are independent of each other. Admittedly, this assumption about the independence of the approximation errors is very hard to verify and may not hold in practice. Nonetheless, the theorem does explain why recalibrating the covariance matrices $P_{y,k},S_{k}$, and $P_{xy,k}$ after the state update is advantageous for accurate state covariance estimation. Specifically, the advantage comes from our replacement of (10) with (18), which highlights the fundamental idea of the new framework. That is, instead of reducing the updated state covariance matrix to the smallest possible value, we re-approximate the system after the state update to estimate the actual effect of the Kalman gain more accurately. In fact, in the new framework, the trace of the state covariance matrix may increase after the state update and recalibration if the system is strongly nonlinear and the predicted states have high variances, just like the example shown in Fig. 1. This phenomenon means that the update fails to make the estimation more accurate and should be withdrawn. To address this situation, we also add a “back out” step to the framework, which allows the state estimator to back out of the update when the update is unhelpful. The pseudo-code of the new framework is shown in Algorithm 1. Note that the new framework degrades to the conventional framework when the “recalibrate” and “back out” steps are replaced with (10).
@@
26:where, $\bm{\hat{x}}$ is the estimated states, $\bm{P}$ is the states’ covariance matrix, $\bm{\hat{y}}$ is the estimated measurements, $\bm{P}_{xy}$ is the covariance between the states and the measurements, $\bm{P}_{y}$ is the covariance of the estimated measurements, $\bm{S}$ is the innovation (or residual) covariance, $\bm{K}$ is the Kalman gain. The subscript $k|k-1$ represents the estimated value before the state update, and $k|k$ represents the estimated value after the state update.
It is worth mentioning that when the measurement functions are all linear, they would require no approximation. Or, to be more precise, the approximation of the measurement functions would become independent of the states. In this case, the approximations of $P_{y,k},S_{k}$, and $P_{xy,k}$ will not change before and after the recalibration, and our framework will be equivalent to the conventional framework.
To give readers an intuitive understanding of the effect of the “recalibrate” and “back out” steps, in Fig. 2, we showed an example of using different types of nonlinear KFs to do the state update given the measurement $y=x^{3}/3-x^{2}/8-x+1.5383=0$. The measurement function is represented by the thick gray curve in the figure. The predicted state is $x_{1|0}=0$, its variance is $1.5^{2}$, and the standard deviation of the measurement is $\sigma_{y}=0.01$. This state estimation problem is challenging because the derivative of the measurement function at the predicted state is negative, while the actual state is to the left of the predicted state. The state and covariance estimations made by different algorithms shown in different rows in the figure. Generally, the conventional framework always underestimates the actual state estimation error, which can be visualized by the narrow standard deviation estimations in Fig. 2 that cannot capture the true state. By contrast, the state covariance estimation made by the new framework is much more accurate. Additionally, the new framework is aware that the state estimation becomes less accurate after the update because the estimated variance increases. Therefore, after the “recalibrate” step, the “back out” step will be triggered, and the new framework will withdraw the update and use the predicted state and the predicted state variance as the final output.
@@
Figure 2: The comparison of the covariance estimation after the state update and recalibration using different frameworks and different types of nonlinear Kalman filters.
@@ 2.4 Relationship to Iterated Kalman Filter
The new framework described above may remind the readers of the iterated Kalman filter (IKF), which also re-approximates the measurement function after calculating the Kalman gain [19, 20, 21]. However, IKF is fundamentally different from the new framework in three ways. Firstly, IKF redoes the approximation and recalculates the Kalman gain iteratively until convergence. However, such a convergence is not guaranteed, so the algorithm is less robust and often requires finetuning the step size of each iteration [22]. Even if the iteration converges, the convergence may take a very long time. By contrast, the new framework requires no iteration, and the re-approximation is only done once. Secondly, IKF was initially only used on EKF, and the deduction of iterated EKF (IEKF) requires an explicit approximation of the measurement functions (rows 9 and 10 in Algorithm 2), which is only possible for EKF and EKF2. Although a few attempts have been made to extend IKF to UKF and CKF [20, 21, 23, 24], a consensus on how to do this robustly has not been reached. Therefore, it is not easy to implement IKF on all types of nonlinear KFs while guaranteeing that the estimator can converge in most cases. Our framework, on the other hand, does not have this problem since there is no convergence issue. Thirdly, and most importantly, the basic idea behind IKF is still to find the optimal Kalman gain and not to estimate the covariance matrix accurately. Namely, IKF always updates the states and covariance matrix simultaneously using (10), resulting in overconfident state estimations. When the algorithm cannot converge properly, the estimation can be even worse because the estimated $P_{y,k},S_{k}$, and $P_{xy,k}$ can be less accurate, which can lead to a higher degree of overconfidence. In general, in terms of making a more accurate estimation, finding the optimal Kalman gain is not as important as making the covariance estimation accurate, as we will show later in the simulation results.
2:Initialization:
3:$\bm{\hat{x}}_{0|0}=\operatorname{\mathbb{E}}[\bm{x}_{0}]$
4:$\bm{P}_{0|0}=\operatorname{\mathbb{E}}[(\bm{\hat{x}}_{0|0}-\bm{x}_{0})(\bm{\hat{x}}_{0|0}-\bm{x}_{0})^{T}]$
5:for every time step $k$ do
6:    Predict:
8:     $\bm{\hat{x}}_{k|k-1}=\bm{f}(\bm{\hat{x}}_{k-1|k-1},\bm{u}_{k-1})$
9:     $\bm{P}_{k|k-1}=\bm{F}_{k}\bm{P}_{k-1|k-1}\bm{F}_{k}^{T}+\bm{Q}_{k}$
10:    Update:
11:     $\bm{x}_{0}=\bm{\hat{x}}_{k|k-1}$
12:     for $i=1$ to 1000 do
13:         $\bm{H}_{k|k-1}=\frac{\partial\bm{h}}{\partial\bm{x}}|_{\bm{x}_{i-1}}$
14:         $\bm{\hat{y}}_{k|k-1}=\bm{h}(\bm{x}_{i-1})+\bm{H}_{k|k-1}(\bm{x}_{i-1}-\bm{\hat{x}}_{k|k-1})$
15:         $\bm{P}_{xy,k|k-1}=\bm{P}_{k|k-1}\bm{H}_{k|k-1}^{T}$
16:         $\bm{P}_{y,k|k-1}=\bm{H}_{k|k-1}\bm{P}_{k|k-1}\bm{H}_{k|k-1}^{T}$
17:         $\bm{S}_{k|k-1,i}=\bm{P}_{y,k|k-1}+\bm{R}_{k}$
18:         $\bm{K}_{i}=\bm{P}_{xy,k|k-1}\bm{S}_{k|k-1,i}^{-1}$
19:         $\bm{x}_{i}=\bm{\hat{x}}_{k|k-1}+\bm{K}_{i}(\bm{z}_{k}-\bm{\hat{y}}_{k|k-1})$
20:         if $i>1$ and $||\bm{x}_{i}-\bm{x}_{i-1}||>||\bm{x}_{i-1}-\bm{x}_{i-2}||$ then
21:             $\bm{S}_{k|k-1,i}=\bm{S}_{k|k-1,i-1}$
22:             $\bm{K}_{i}=\bm{K}_{i-1}$
23:             $\bm{x}_{i}=\bm{x}_{i-1}$
24:             break
25:         end if
26:         if $\max_{j}|1-\bm{x}_{i}(j)/\bm{x}_{i-1}(j)|<0.001$ then
27:              break
28:         end if
29:     end for
30:     $\bm{S}_{k|k-1}=\bm{S}_{k|k-1,i}$
31:     $\bm{K}=\bm{K}_{i}$
32:     $\bm{\hat{x}}_{k|k}=\bm{x}_{i}$
33:     $\bm{P}_{k|k}=\bm{P}_{k|k-1}-\bm{K}\bm{S}_{k|k-1}\bm{K}^{T}$
34:end for
35:Note that rows 16–21 are added to prevent the state estimations from diverging. Namely, when the sign of divergence is detected (row 16), the newest updates of the states, Kalman gain, and measurement residual covariance matrix are withdrawn (rows 17–19). If such a withdrawal is not made, we found in the simulation that the estimation accuracy can be much lower. This characteristic also suggests that the algorithm is not very robust and stable.
@@
Since IKF also uses (10) to update the state covariance matrix, it can also be considered as another framework for nonlinear KFs that is similar to the conventional framework. A natural question is: can we combine the new framework with the IKF to get a higher estimation accuracy? Unfortunately, the answer is no because a prerequisite for the new framework to give accurate covariance estimation is that the approximations made at the predicted states (${\color[rgb]{1,0,0}P_{y,k|k-1}},{\color[rgb]{1,0,0}S_{k|k-1}}$, and ${\color[rgb]{1,0,0}P_{xy,k|k-1}}$) are independent of the approximations made at the updated states (${\color[rgb]{1,0,0}P_{y,k|k}},{\color[rgb]{1,0,0}S_{k|k}}$, and ${\color[rgb]{1,0,0}P_{xy,k|k}}$). However, after IKF converges, these two sets of approximations become the same, which violates this prerequisite in Theorem 2. For example, the IEKF algorithm shown in Algorithm 2 approximates the measurement function iteratively at $x_{i-1}$ (rows 9–13). When the algorithm converges, in the final iteration, we have $x_{i-1}\approx x_{i}=\hat{x}_{k|k}$, meaning that the approximation of the measurement functions in the final iteration will stay the same after the state update. In fact, to the best of our knowledge, IKF-related algorithms are the only existing types of nonlinear KFs that can not get large benefits from the proposed framework. Nevertheless, for any type of nonlinear KF that can be combined with IKF, its accuracy can be greatly improved when it is combined with our new framework instead. In this sense, we can still say that all existing types of nonlinear KFs can benefit from the proposed framework.

### v11

@@ 2.3 The proposed new framework for nonlinear Kalman filters
Note that the proposed new framework selects the same Kalman gain and does the same state update as the conventional one; the only difference is that the state covariance update equation is now a direct approximation of (14), which is the actual state covariance regardless of the selection of the Kalman gain $K$. Intuitively, this can prevent the KF from overconfidence because the equation no longer relies on the problematic assumption that the Kalman gain is optimal. To explain this advantage more rigorously, we consider the newly approximated covariance matrices ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{z,k|k}},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}S_{k|k}}$, and $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{xz,k|k}$} produced by a nonlinear KF as (random) estimators of the actual quantities $P_{z,k}$, $S_{k}$, and $P_{xz,k}$. Similar to our previous assumptions about ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{z,k|k-1}},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}S_{k|k-1}}$, and $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{xz,k|k-1}$}, we regard $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{xz,k|k}$} as a random matrix, $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{z,k|k}$} as a random positive semidefinite matrix, and $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}S_{k|k}$} as a random positive definite matrix. Additionally, we assume that these approximated covariance matrices are unbiased estimators and satisfy:
@@
$\begin{cases}\operatorname{\mathbb{E}}[{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{z,k|k}}]=P_{z,k}\\ \operatorname{\mathbb{E}}[{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}S_{k|k}}]=S_{k}\\ \operatorname{\mathbb{E}}[{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{xz,k|k}}]=P_{xz,k}\end{cases}$ (28)
@@
Assume that the conditions (18) and (28) hold. Further assume that the random pair $({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{xz,k|k-1}},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}S_{k|k-1}})$ is independent of the random pair $({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{xz,k|k}},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}S_{k|k}})$. Then $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\Delta P_{k|k,\mathrm{err,new}}$}, as defined in (27), satisfies
$\operatorname{\mathbb{E}}\big[{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\Delta P_{k|k,\mathrm{err,new}}}\big]=0.$ (29)
Consequently, if in addition $\operatorname{\mathbb{E}}[{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{k|k-1,\text{est}}}]=P_{k|k-1}$, then $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{k|k,\mathrm{est,new}}$} calculated in (25) is an unbiased estimator of $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{k|k,\mathrm{ac}}$}. Namely,
$\operatorname{\mathbb{E}}\big[{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{k|k,\mathrm{est,new}}}-{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{k|k,\mathrm{ac}}}\big]=0.$ (30)
@@
$\operatorname{\mathbb{E}}[{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\Delta P_{k|k,\mathrm{err,new}}}]$ (31)
$=$ $\operatorname{\mathbb{E}}[{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{xz,k|k-1}}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}S_{k|k-1}^{-1}}({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}S_{k|k}}-S_{k}){\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}S_{k|k-1}^{-1}}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{xz,k|k-1}^{T}}]$
$-\operatorname{\mathbb{E}}[({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{xz,k|k}}-P_{xz,k}){\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}S_{k|k-1}^{-1}}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{xz,k|k-1}^{T}}]$
$-\operatorname{\mathbb{E}}[{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{xz,k|k-1}}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}S_{k|k-1}^{-1}}({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{xz,k|k}}-P_{xz,k})^{T}]$
$=$  $0.$
@@
The last equality holds because the terms in parentheses are zero, due to the unbiasedness (i.e., $\operatorname{\mathbb{E}}({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}S_{k|k}})=S_{k}$ and $\operatorname{\mathbb{E}}({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{xz,k|k}})=P_{xz,k}$) and independence assumptions.
@@
Theorem 2 suggests that the proposed new framework can generally prevent the filter from being overconfident in the “update” step if the approximation errors of the measurement function at the predicted and updated states are independent of each other. Admittedly, this assumption about the independence of the approximation errors is very hard to verify and may not hold in practice. Nonetheless, the theorem does explain why recalibrating the covariance matrices $P_{z,k},S_{k}$, and $P_{xz,k}$ after the state update is advantageous for accurate state covariance estimation. Specifically, the advantage comes from our replacement of (13) with (25), which highlights the fundamental idea of the new framework. That is, instead of reducing the updated state covariance matrix to the smallest possible value, we re-approximate the system after the state update to estimate the actual effect of the Kalman gain more accurately.
In nonlinear KFs, the update can sometimes make the state estimate less accurate and increase the actual state covariance, as shown in Fig. 1. If this situation were detectable, a natural reaction would be to withdraw the update and use the more accurate predicted state (and the predicted state covariance) as the final output. Unfortunately, such detection is impossible based solely on the estimated state covariance matrix in the conventional KF framework, since $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{k|k,\mathrm{est}}$} calculated in (13) is guaranteed to be smaller than or equal to $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{k|k-1,\mathrm{est}}$}. In the new framework, however, the additional “recalibrate” step makes such detection possible, which motivates us to introduce a “back out” step. In general, “back out” means to withdraw unhelpful updates and output the state estimate with a lower estimated mean-squared error. Specifically, by the end of the recalibrate step, the filter has two candidate estimates of the current state: the predicted state $\hat{x}_{k|k-1}$ with covariance $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{k|k-1,\mathrm{est}}$}, and the updated state $\hat{x}_{k|k}$ with covariance $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{k|k,\mathrm{est,new}}$}. Since the trace of the covariance matrix equals the expected squared Euclidean error of the corresponding state estimate, we choose as the final output the estimate whose covariance has the smaller trace. If the “back out” step is omitted, the filter may not only produce a suboptimal estimate, but can also become highly unstable: $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{k|k,\mathrm{est,new}}$} computed in (25) is no longer guaranteed to remain bounded by $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{k|k-1,\mathrm{est}}$}, and can be several orders of magnitude larger than $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{k|k-1,\mathrm{est}}$}, which is clearly undesirable.
The pseudo-code of the new framework is shown in Algorithm 1. Note that the new framework degrades to the conventional framework when the “recalibrate” and “back out” steps are replaced with (13). Moreover, when all measurement functions are linear, the approximations of $P_{z,k}$, $S_{k}$, and $P_{xz,k}$ become identical before and after the recalibration step. In this case, the proposed framework is exactly equivalent to the conventional framework and preserves its optimality for linear systems.
@@
26:where, $\bm{\hat{x}}$ is the estimated states, $\bm{P}$ is the states’ covariance matrix, $\bm{\hat{z}}$ is the estimated measurements, $\bm{P}_{xz}$ is the covariance between the states and the measurements, $\bm{P}_{z}$ is the covariance of the estimated measurements, $\bm{S}$ is the innovation (or residual) covariance, $\bm{K}_{k}$ is the Kalman gain. The subscript k|k1k|k-1 represents the estimated value before the state update, and $k|k$ represents the estimated value after the state update.
To provide an intuitive illustration of the roles of the “recalibrate” and “back out” steps, Fig. 2 shows a single update for several nonlinear KFs applied to the scalar measurement model
$z=h(x)+v,\qquad h(x)=x^{3}/3-x^{2}/8-x+1.5383,$ (32)
with the observed measurement $z_{1}=0$. The measurement function $h(x)$ is plotted as the thick gray curve. The predicted state is $x_{1|0}=0$ with variance $1.5^{2}$, and the measurement noise standard deviation is $\sigma_{y}=0.01$. This example is particularly challenging because $h^{\prime}(x_{1|0})<0$ while the true state lies to the left of $x_{1|0}$, so a local linearization can easily mislead the filter. The state and covariance estimates produced by different algorithms are shown in separate rows in the figure. Under the conventional framework, all methods substantially underestimate the actual estimation error, as indicated by the overly narrow $\pm 1\sigma$ intervals that fail to cover the true state. In contrast, within the proposed framework, the covariance estimates are much more consistent with the true error. In this example, the recalibrated variance correctly indicates that the update is less reliable than the prediction, so the “back out” step is triggered and the filter discards the misleading update, outputting the predicted state and its variance as the final estimate. If the “back out” step is removed from the framework, the estimated state becomes less accurate for all filters. Moreover, in this example, the CKF’s estimated variance after the recalibration is approximately $144^{2}$, which is almost 10,000 times larger than the variance of the predicted state. Such a “covariance explosion” further underscores the necessity of the “back out” step.
@@
Figure 2: The comparison of the variance estimation after the state update and recalibration using different frameworks and different types of nonlinear Kalman filters. The recalibrated variance of CKF (new) is approximately $144^{2}$, almost 10,000 times larger than the predicted variance, which underscores the importance of the “back out” step.
Finally, it is insightful to interpret the proposed framework through a machine learning perspective. In the conventional nonlinear KF framework, the Kalman gain is chosen based on approximations constructed around the predicted state. These local approximations can be viewed as an analogue of a “training set”: they are reasonably representative of the underlying nonlinear system but do not capture its full complexity. Computing the covariance using the same approximations is then analogous to evaluating an in-sample (training) loss, which tends to be systematically smaller than the true out-of-sample (testing) loss. This offers an intuitive explanation for why the conventional nonlinear KF framework often yields overconfident covariance estimates. In contrast, our new framework computes the covariance based on approximations constructed around the updated state, which can be viewed as an analogue of a “validation set.” In this interpretation, Theorem 2 states that, when the information used to design the gain and the information used to assess the covariance are independent, the resulting covariance (or loss) estimate is unbiased. Even though exact independence may not hold in practice, the machine learning literature provides strong empirical evidence that incorporating a validation-like mechanism is effective in mitigating overconfidence (overfitting) and improving estimation accuracy, which is precisely the role played by the “recalibrate” step introduced in this paper. As for the “back out” step, it can be interpreted as reverting to the previous checkpoint when the validation error increases during training, which is also necessary and important for finding the best estimates.
@@ 2.4 Relationship to Iterated Kalman Filter
The new framework described above may remind the readers of the iterated Kalman filter (IKF), which also re-approximates the measurement function after calculating the Kalman gain [29, 30, 31]. However, IKF is fundamentally different from the new framework in three ways. Firstly, IKF redoes the approximation and recalculates the Kalman gain iteratively until convergence. However, such a convergence is not guaranteed, so the algorithm is less robust and often requires finetuning the step size of each iteration [32]. Even if the iteration converges, the convergence may take a very long time. By contrast, the new framework requires no iteration, and the re-approximation is only done once. Secondly, IKF was initially only used on EKF, and the deduction of iterated EKF (IEKF) requires an explicit approximation of the measurement functions (rows 9 and 10 in Algorithm 3 in Appendix C), which is only possible for EKF and EKF2. Although a few attempts have been made to extend IKF to UKF and CKF [30, 31, 33, 34], a consensus on how to do this robustly has not been reached. Therefore, it is not easy to implement IKF on all types of nonlinear KFs while guaranteeing that the estimator can converge in most cases. Our framework, on the other hand, does not have this problem since there is no convergence issue. Thirdly, and most importantly, the basic idea behind IKF is still to find the optimal Kalman gain and not to estimate the covariance matrix accurately. Namely, IKF always updates the states and covariance matrix simultaneously using (13), resulting in overconfident state estimations. When the algorithm cannot converge properly, the estimation can be even worse because the estimated $P_{z,k},S_{k}$, and $P_{xz,k}$ can be less accurate, which can lead to a higher degree of overconfidence. In general, in terms of making a more accurate estimation, finding the optimal Kalman gain is not as important as making the covariance estimation accurate, as we will show later in the simulation results.
@@
Since IKF also uses (13) to update the state covariance matrix, it can also be considered as another framework for nonlinear KFs that is similar to the conventional framework. A natural question is: can we combine the new framework with the IKF to get a higher estimation accuracy? Unfortunately, the answer is no because a prerequisite for the new framework to give accurate covariance estimation is that the approximations made at the predicted states (${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{z,k|k-1}},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}S_{k|k-1}}$, and $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{xz,k|k-1}$}) are independent of the approximations made at the updated states (${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{z,k|k}},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}S_{k|k}}$, and $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{xz,k|k}$}). However, after IKF converges, these two sets of approximations become the same, which violates this prerequisite in Theorem 2. For example, the IEKF algorithm shown in Algorithm 3 approximates the measurement function iteratively at $x_{i-1}$ (rows 9–13). When the algorithm converges, in the final iteration, we have $x_{i-1}\approx x_{i}=\hat{x}_{k|k}$, meaning that the approximation of the measurement functions in the final iteration will stay the same after the state update. In fact, to the best of our knowledge, IKF-related algorithms are the only existing types of nonlinear KFs that can not get large benefits from the proposed framework. Nevertheless, for any type of nonlinear KF that can be combined with IKF, its accuracy can be greatly improved when it is combined with our new framework instead. In this sense, we can still say that all existing types of nonlinear KFs can benefit from the proposed framework.